# Lossy Compression with Data, Perception, and Classification Constraints

### v1

@@ III-B Gaussian Source
For a scalar Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the information rate-distortion-classification function with square error distortion is given by
@@
We also visualize the result of Theorem 2 by plotting the $R(D,C)$ function of a Gaussian source as a surface as well as distortion-classification curves for different rates (Fig. 2(b)). Similar properties with binary source such as convexity and non-increasing of $R(D,C)$ on $D$ and $C$ could be observed.
Figure 2: Visualization for RDC function of (a) a Bernoulli source and (b) a Gaussian source. The left figures depict $R(D,C)$ function as a surface. The colored solid lines on the figure indicate the different rate levels. The right figures show $R(D,C)$ function along distortion-classification planes, which is the two-dimensional projection of contour lines of (a).
In addition, we investigate the activeness of distortion constraint and classification constraint given different value pairs of $(D,C)$. As illustrated in Fig. 3, $D$ and $C$ are at odds with each other in an ‘antagonistic’ way where once $D$ is relatively small ($D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{2C-h(S)}))$) and the distortion constraint is active, the classification constraint would be silent, resulting in the dominance of $D$ in determining the rate. Conversely, as $D$ increases, the classification will be the only active constraint and the rate is determined only by $C$.
@@
The phenomenon differs from the case of RDP [8][9]. As proved in [9], in Gaussian case with Wasserstein-2 perception loss, when $\sqrt{P}\geq\sigma_{X}-\sqrt{|\sigma_{x}^{2}-D|}$, the distortion constraint becomes the only active constraint, and the rate is solely determined by $D$. However, as $P$ decreases, both $P$ and $D$ are active, leading to a joint determination of the rate by both $D$ and $P$. Note that $I(X,\hat{X})=0$ is achievable for any $P$ when $D$ is not an active constraint[8].
IV Rate-perception-classification Relationship
In practice, generative tasks and classification tasks are often jointly considered in the context of conditional generative adversarial nets (GAN)[25], with great potential in controlling the categories of generated images. In this section, we aim to investigate the inherent relationship between rate-perception-classification (RPC).
Specifically, let us consider the following information rate-perception-classification (RPC) function where distortion constraint (1a) is relaxed in RDPC function and only one classification variable is considered:
@@ IV Rate-perception-classification Relationship
IV-A Binary source
We now find the optimal rate $R(P,C)$ required to describe a binary source with perception quality characterized by total variation divergence.
@@ IV-A Binary source
For a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$), if $C<H(S_{1})$, the problem is infeasible; otherwise, the information rate-distortion-classification function with total variation (TV) divergence is given by
@@
The conditional entropy constraint provides a lower bound on the rate. By delicately assigning the value of conditional probability $p_{\hat{X}|X}(\hat{X}=0|X=0)$ and $p_{\hat{X}|X}(\hat{X}=0|X\!=\!1)$, the lower bound is achievable with the TV divergence $d_{TV}(p_{X},p_{\hat{X}})\!=\!0$. See the details in Appendix C. ∎
The above theorem indicates that the rate only depends on the classification constraint, and there is no tradeoff between classification and perception. The proof of Theorem 3 establishes that it is always possible to find an optimal solution that achieves a TV divergence of zero while attaining the lower bound of the rate imposed by the classification constraint.
IV-B Gaussian source
The similar conclusion could be obtained for the Gaussian case, where the optimal rate is solely determined by the level of the conditional entropy constraint, and the perception constraint is always inactive. In this subsection, we take the natural base of logarithm.
@@ IV-B Gaussian source
For a scalar Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the information rate-perception-classification function with KL divergence is given by
@@
We will prove that the lower bound given by the conditional entropy constraint is achievable with the KL divergence always being zero. See the details in Appendix D. ∎
@@ V Experimental Results
In this section, we verify our theoretical results by implementing a deep-learning (DL) based image compression framework incorporating different constraints. We illustrate the characteristics of RDC and RPC tradeoffs in the MNIST dataset, which coincide with our theoretical results.
@@
Similar to [9], we utilize a generative adversarial network (GAN) with stochastic encoder and decoder. As shown in Fig. 4, the system consists of an encoder $E$, a decoder $G$, a discriminator $D$, and a pre-trained classifier $C$.
The output reconstruction $\hat{x}$, as well as the original image $x$, are fed to the discriminator $G$ and compute the discriminator loss. The reconstruction $\hat{x}$ is also fed to a pre-trained classifier to obtain a one-hot vector $\hat{\mathbf{s}}$ with each entry being the probability that the image belongs to the corresponding class. The true label $s$ and the predicted $\mathbf{\hat{s}}$ are used to compute the classification accuracy.

### v2

@@ III-A2 Gaussian source
Consider a Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$. The problem is infeasible if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$. Otherwise, the information rate-distortion-classification function with MSE distortion is given by
@@
We also visualize the result of Theorem 2 by plotting the $R(D,C)$ function of a Gaussian source as a surface (Fig. 2(b), left) as well as distortion-classification curves for different rates (Fig. 2(b), right). Similar properties with binary case such as convexity and non-increasing of $R(D,C)$ on $D$ and $C$ are also observed.
Figure 2: Visualization for RDC function of (a) a Bernoulli source and (b) Gaussian source. The left figures depict $R(D,C)$ function as a surface. The colored solid lines on the figure indicate the different rate levels. The right figures show $R(D,C)$ function along distortion-classification planes, which is the two-dimensional projection of contour lines of the left figures.
In addition, we investigate the activeness of distortion constraint (2a) and classification constraint (2b) given different value pairs of $(D,C)$. As depicted in Fig. 3, there is an ’antagonistic’ relationship between D and C. When D is relatively small ($D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))$) and the distortion constraint is active, the classification constraint becomes inactive. Consequently, D predominates in determining the rate. Conversely, as $D$ increases, the classification will be the only active constraint and the rate is determined by $C$.
@@
The phenomenon differs from the case of RDP [8, 11, 34]. As proved in [11], in the Gaussian case, when $\sqrt{P}\geq\sigma_{X}-\sqrt{|\sigma_{x}^{2}-D|}$, the distortion constraint becomes the only active constraint, and the rate is solely determined by $D$. However, as $P$ decreases, both $P$ and $D$ are active, leading to a joint determination of the rate by $D$ and $P$. Here $D$ is always active since $I(X,\hat{X})=0$ is achievable for any $P$ when $D$ is not an active constraint[8]. Recently, a more general diagram of the transition curves determining the activeness of $P$ and $D$ in the RDP function has been depicted in [34].
III-B Rate-perception-classification Relationship
Unlike the cases of RDP and RDC, we will show that there is no tradeoff between perception and classification in the RPC relationship for both binary and Gaussian scenarios. Intuitively, the restriction on target label inference may not affect the potential of perfect perceptual quality.
Specifically, let us consider the following information rate-perception-classification function where the distortion constraint (1a) is relaxed in the RDPC function and only one classification variable is considered:
@@ III-B Rate-perception-classification Relationship
III-B1 Binary source
The following theorem characterizes $R(P,C)$ for a binary source with perception quality described by total variation divergence.
@@ III-B1 Binary source
Consider a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). If $C<H(S_{1})$, the problem is infeasible. Otherwise, the information rate-distortion-classification function with total variation (TV) divergence is given by
@@
The conditional entropy constraint provides a lower bound on the rate. By delicately assigning the value of conditional probability $P(\hat{X}=0|X=0)$ and $P(\hat{X}=0|X=1)$, the lower bound is achievable with the TV divergence $d_{TV}(p_{X},p_{\hat{X}})=0$. See details in Appendix C. ∎
The above theorem indicates that there is no tradeoff between perception and classification in the RPC relationship, since the rate only depends on the classification constraint. The proof of Theorem 3 establishes that it is always possible to find an optimal solution that achieves a TV divergence of zero while attaining the lower bound of the rate imposed by the classification constraint.
III-B2 Gaussian source
The conclusions in the binary case can be extended to the Gaussian case, where the optimal rate is solely determined by the level of the conditional entropy constraint, and the perception constraint is always inactive. In this subsection, we take the natural base of logarithm.
@@ III-B2 Gaussian source
Consider a scalar Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$. If $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible. Otherwise, the rate-perception-classification function with KL divergence is given by
@@
In Appendix D, we will prove that the lower bound given by the conditional entropy constraint is achievable with the KL divergence always zero. ∎
@@
The proposed RDC and RPC functions complement the IB principle and RDP tradeoff, providing new insights regarding the interplay between different tasks under a limited rate. Specifically, the perceptual quality primarily relates to the distribution of the data, which does not inherently create a tradeoff with the rate. However, when constrained by a limited rate, the distortion and classification constraints have distinct impacts on the perceptual quality. That is, minimizing distortion potentially leads to degradation of perceptual quality, while imposing a conditional entropy constraint does not limit the ability to perfectly preserve the distribution, at least in the binary and Gaussian scenarios.
On the other hand, when the rate is not limited, it becomes possible to reconstruct the source information perfectly. So, there should be no tradeoffs between the different tasks if the rate is sufficiently large. However, within the framework of signal restoration, previous literature[21] has demonstrated the existence of a classification-distortion-perception (CDP) tradeoff without an explicit constraint on the communication rate. In the next section, we will initiate a discussion to identify the critical factor that determines the presence of tradeoffs.
IV The Decisive Role of Source Noise in Tradeoffs
In this section, we will first compare the models of lossy compression and signal restoration. For the latter, an extrinsic noise is first introduced to degrade the source $X$ and the goal is to restore $\hat{X}$ from the degraded observation $Y$. With a toy example, we will illustrate that in the scenario of non-zero noise level, the tradeoffs between distortion, perception and classification emerge. Then, we show that introducing an extrinsic noise is equivalent to imposing a certain level of distortion between $X$ and $\hat{X}$. With this equivalence, we further consider the RPC relationship given a specific level of distortion and observe different characteristics compared to the original RPC tradeoff. The practical role of these two tradeoffs will be verified by the experiments in Section V.
IV-A Comparison of Lossy Compression and Signal Restoration
Figure 4: Frameworks of lossy compression model and signal restoration model.
Fig. 4 illustrates a comparison between the models of the lossy compression and the signal restoration. In the lossy compression setting, our objective is to find an efficient representation $f(X)$ that minimizes the communication bits given a constraint of distortion (or perception/classification in the task-oriented model). From an information-theoretic perspective, we directly optimize the conditional distribution $p_{\hat{X}|X}$ by jointly designing the encoder and decoder. In the signal degradation and restoration model, an extrinsic noise is first introduced through $p_{Y|X}$ to the source $X$, resulting in the observation of a degraded source $Y$. The task is to reconstruct $\hat{X}$ from the degraded $Y$. In the following, we mainly consider the additive Gaussian noise, i.e., $Y=X+N$ with $N\sim\mathcal{N}(0,\sigma_{N}^{2})$. To delve deeper into the tradeoffs in the signal restoration model, let us begin by considering the toy example in [21].
Example 1 (Toy example in [21]).
Consider signal $X$ with Gaussian mixture distribution $p_{X}(x)=P_{1}p_{X_{1}}(x)+P_{2}p_{X_{2}}(x)$, where $P_{1}=0.7,P_{2}=0.3$, $p_{X_{1}}(x)=\mathcal{N}(-1,1)$, $p_{X_{2}}(x)=\mathcal{N}(1,1)$. The signal degradation is given by $Y=X+N$ where $N\sim\mathcal{N}(0,\sigma_{N}^{2})$ with $\sigma_{N}=1$. Consider the linear denoising method $\hat{X}=aY$ where $a$ is an adjustable parameter.
According to the Bayes decision rule, the classification error rate given by the optimal classification plane $c_{0}$ is
[…]
$=P_{2}\Phi(x_{0}^{\prime})+P_{1}\Phi(-x_{0}^{\prime\prime}),$
where $x_{0}^{\prime}=\frac{x_{0}-a}{|a|\sqrt{1+\sigma_{N}}}$ and $x_{0}^{\prime\prime}=\frac{x_{0}+a}{|a|\sqrt{1+\sigma_{N}}}$, and $\Phi(\cdot)$ is the cumulative distribution function (CDF) of the standard normal distribution. Note that the error rate is thus a function of $a$.
Meanwhile, the KL divergence $d_{KL}(p_{X},p_{\hat{X}})$ and MSE $\mathbb{E}[\Delta(X,\hat{X})]$ are also functions of the denoising parameter $a$ which can be computed numerically.
Specifically, MSE as a function of $a$ is plotted in Fig. 5. Note that the optimal MSE is $0.204$ obtained at $a=0.67$. No matter how we choose the value of $a$, MSE will never be 0, which means perfect reconstruction is impossible.
Figure 5: MSE as functions of linear denoising parameter $a$ in the toy example.
Meanwhile, all three relationships, namely distortion-perception (DP), distortion-classification (DC), and perception-classification (PC) exhibit tradeoffs, as shown in Fig. 6. For example, in the case of the DP tradeoff shown in Fig. 6(a), we draw the pairs $(D,P)$ as the solution of
$P=$ $\min_{a}\ d_{KL}(p_{X},p_{\hat{X}})$
[…]
and it shows that the decrease in MSE is accompanied by an increase in KL divergence.
Figure 6: Distortion-perception tradeoff (a), distortion-classification tradeoff (b) and perception-classification tradeoff (c) in the toy example.
Figure 7: Simulation of rate-perception-classification function given a certain level of distortion $R(P,C|D)$ in the Gaussian case where $\sigma_{x}=1,\sigma_{s}=0.7,\theta_{1}=0.63$. (a) The given distortion $D=0.5$; (b) The given distortion $D=0.6$; (c) The given distortion $D=0.8$.
Similar to the case of the sufficiently large rate in the lossy compression framework, when the noise level $\sigma_{N}$ tends to zero here, there are no tradeoffs between different tasks. In this scenario, we directly observe the original data. The identity restoration by setting $a=1$ yield with zero MSE, zero KL divergence, and the optimal classification performance.
However, when dealing with a non-zero noise level, we observe that the objectives of minimizing MSE, error rate, and KL divergence do not always align, leading to tradeoffs between each pair of constraints. Specifically, for a linear denoiser, the optimal MSE is achieved at $a=0.67$, the minimum error rate is achieved at $a=0.50$, and the lowest KL-divergence is obtained at $a=0.81$ for $\sigma_{N}=1$. The value of optimal parameter $a$ depends on the noise level $\sigma_{N}$.
In the following, we show that introducing an non-zero extrinsic noise is equivalent to imposing a certain level of distortion between $X$ and $\hat{X}$. Starting with a simple example, where the degradation is given by $Y=X+N\sim\mathcal{N}(0,1+\sigma_{N}^{2})$ and the restoration is given by $\hat{X}=aY\sim\mathcal{N}(0,a^{2}(1+\sigma_{N}^{2}))$, for distortion measured in MSE, we have
$\mathbb{E}[\Delta(X,\hat{X})]$ $=1+a^{2}(1+\sigma_{N}^{2})-2a$
$=(1+\sigma_{N}^{2})(a-\frac{1}{1+\sigma_{N}^{2}})^{2}+\frac{\sigma_{N}^{2}}{1+\sigma_{N}^{2}},$
which is always greater than zero as long as $\sigma_{N}\neq 0$. In general, based on the results from lossy compression with noisy sources [35, 36], when the distortion is measured by MSE, we have
$\mathbb{E}[||X-\hat{X}||^{2}]=\mathbb{E}[||X-\mathbb{E}[X|Y]||^{2}]+\mathbb{E}[||\hat{X}-\mathbb{E}[X|Y]||^{2}],$
where the term $\mathbb{E}[||X-\mathbb{E}[X|Y]||^{2}]$ determines the minimum distortion, which is independent of the encoder-decoder pair.
Thus, the presence of source noise introduces an implicit constraint on distortion, making perfect reconstruction impossible. Detailed investigation of the tradeoffs inherent in lossy compression with a noisy source remains a topic for future research. In the following, we verify the significant role of source noise in the existence of tradeoffs from the perspective of imposing equivalent distortion constraints. Specifically, we reexamine the RPC relationship in Section III-B with a specific level of distortion. Through this analysis, we observe that the tradeoff between perception and classification indeed exists.
IV-B RPC Given a Certain Level of Distortion
In section III-B, we found that there is no tradeoff between perception and classification in the RPC relationship The optimal rate only depends on the classification constraint. However, our subsequent simulations in the scalar Gaussian case will demonstrate that imposing a constraint on the distortion level causes a tradeoff to emerge.
Without loss of optimality, the RPC function for a scalar Gaussian case is equivalent to (see Appendix D)
$R(P,C)=$ $\min_{\sigma_{\hat{x},}\theta_{2}}-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$
s.t. $\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}}\leq P,$
$-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})\geq h(S)-C,$
where $\sigma^{2}_{\hat{X}}$, $\theta_{1}$ and $\theta_{2}$ denote the variance of $\hat{X}$, $Cov(S,X)$ and $Cov(\hat{X},X)$ respectively. If we consider a specific level of $D$, i.e., $\mathbb{E}[||X-\hat{X}||^{2}]=\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-2\theta_{2}=D$, by substituting $\theta_{2}=\frac{\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-D}{2}$ to the RPC function, we have
$R(P,C|D)=\min_{\sigma_{\hat{x}}}$ $-\frac{1}{2}\log(1-\frac{(\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-D)^{2}}{4\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$
s.t. $\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}}\leq P,$
$\!\!\!\!\!\!\!\!\!\!\!-\!\frac{1}{2}\log(1\!-\!\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{(\sigma_{x}^{2}\!+\!\sigma_{\hat{x}}^{2}\!-\!D)^{2}}{4\sigma_{\hat{x}}^{2}})\!\geq\!h(S)\!-\!C.$
Since the objective and two constraints are all functions of $\sigma_{\hat{x}}$, we can depict the curve of $R(P,C|D)$ by simulation, as shown in Fig. 7. Here, we choose $\sigma_{x}=1,\sigma_{s}=0.7,\theta_{1}=0.63$ and $D=0.5,0.88$ respectively.
It is observed that the perception and classification exhibit a tradeoff, which shrinks as $D$ decreases. This can be explained by considering the interplay of $D$ with classification and perception respectively. From the Gaussian RDP tradeoff[11], we know that when $P>\sigma_{x}-\sqrt{|\sigma_{x}^{2}-D|}$, the perception constraint is inactive and the rate only depends on $D$. Meanwhile, in Section III-A2, we have shown that when $C\geq\frac{1}{2}\log(1-\rho^{2}(1-\frac{D}{\sigma_{x}^{2}}))+h(S)$, the rate only depends on $D$. Hence, when $D\rightarrow 0$, a perfect reconstruction is expected, and the rate is dominated by $D$. When $D$ is a bit larger and a degraded reconstruction is expected, the objectives of optimizing perception and classification are not always aligned and the rate will depend on the activeness of $C$ or $P$ constraints. Thus, we can see that the tradeoff emerges.
@@ V Experimental Results
In this section, we validate our theoretical results by implementing a DL-based image compression framework to achieve multiple objectives related to distortion, perception, and classification. These experiments not only verify the effectiveness of our theoretical results but also provide insights into loss function design in multi-task learning.
@@
V-A DL-based Lossy Compression with RDPC Constraints
By utilizing a generative adversarial network (GAN), the process of loss compression $p_{\hat{X}|X}(\hat{x}|x)$ can be achieved with a encoder and a decoder parameterized by adjustable parameter $\phi_{e}$ and $\phi_{d}$ respectively. Let $\phi=\{\phi_{e},\phi_{d}\}$ denote the parameters of the whole encoder-decoder network.
Similar to [11], we use a GAN-based network with stochastic encoder and decoder with universal quantization[37, 38]. As shown in Fig. 8, the system consists of an encoder $E$, a decoder $G$, a discriminator $D$, and a pre-trained classifier $C$.
The reconstruction $\hat{x}$ and the original image $x$ are fed to the discriminator $G$ to compute the discriminator loss. The reconstruction $\hat{x}$ is also fed to a pre-trained classifier to obtain a one-hot vector $\hat{\mathbf{s}}$ with each entry being the probability that the image belongs to each class. The true label $s$ and the predicted $\mathbf{\hat{s}}$ are used to compute the classification accuracy.
To compute the conditional entropy constraints, we need to track the posterior distribution $p_{\phi}(s|\hat{x})$, which can be replaced by the pre-trained classifier parameterized by $\psi$. By introducing the approximation $p_{\psi}(s|\hat{x})$, we can derive an upper bound of the conditional entropy constraint[39]:
$H(S|\hat{X})\!$ $=\!\!\sum_{s}\sum_{\hat{x}}p_{\phi}(s,\hat{x})\log\frac{1}{p_{\phi}(s|\hat{x})}$
$\!\!\!\!\!\!\!\!\!\!=\!\!\underbrace{\sum_{s}\!\sum_{\hat{x}}\!p_{\phi}(s,\!\hat{x})\!\log\!\frac{p_{\psi}(s|\hat{x})}{p_{\phi}(s|\hat{x})}}_{-d_{KL}(p_{\phi}(s|\hat{x})||p_{\psi}(s|\hat{x}))}\!+\!\!\underbrace{\sum_{s}\!\sum_{\hat{x}}p_{\phi}(s,\hat{x})\!\log\!\frac{1}{p_{\psi}(s|\hat{x})}}_{\text{CE}(s,\mathbf{\hat{s}})}$
$\!\!\!\!\!\!\!\!\!\!\leq\text{CE}(s,\mathbf{\hat{s}}),$
where the inequality follows from the non-negativity of KL- divergence, and $\text{CE}(s,\mathbf{\hat{s}})$ is the cross entropy loss.
For the reconstruction constraint, we use MSE $\mathbb{E}(||X-\hat{X}||^{2})$ serving as the traditional distortion loss. For the perception constraint, we employ the Wasserstein-1 loss to measure the distance between distributions.